OpenStudy (anonymous):

10 - 2(5) -------- 4 -5 Can someone please help solve this fraction

OpenStudy (anonymous):

0

OpenStudy (anonymous):

how did you get that answer

OpenStudy (anonymous):

this a fraction problem 10 -2(5) top, bottom 4 - 5

OpenStudy (anonymous):

10-2(5) = 10 - 10 = 0 ------- ------- --- 4-5 = -1 = -1 0 over anything is 0 :)